Zorna weighs 91 lb. Her weight is 6lb more than half on her fathers weight.how much does her father weigh?
Answer by checkley71(8403)

91=X/2+6
X/2=91-6
X/2=85
X=170 LB. IS THE WEIGHT OF THE FATHER.
PROOF:
91=170/2+6
91=85+6
91=91